OpenStudy (anonymous):

Express 9^x in terms of y, where y=3^x. Don't precisely know what they want here?

OpenStudy (anonymous):

Or, rather, I don't precisely know what the are asking.

OpenStudy (anonymous):

I got y = sqrt(3^2x) which is just y = 3x

OpenStudy (anonymous):

or y^2 = 9^x

OpenStudy (anonymous):

\[y=3^x~~\Rightarrow~~y^2=(3^x)^2=3^{2x}=(3^2)^x=9^x~~\Rightarrow~~y=\pm\sqrt{9^x}\]

OpenStudy (anonymous):

Righto thanks!